# Multi-Preference Lambda-weighted Listwise DPO for Dynamic Preference Alignment

### v4

@@ Abstract
While large language models (LLMs) excel at text generation, aligning them with human preferences remains challenging. Reinforcement learning from human feedback (RLHF) improves alignment but is costly and unstable. Direct Preference Optimization (DPO) offers a simpler alternative, yet assumes a fixed, single-dimensional preference.
We propose Multi-Preference Lambda-weighted Listwise DPO, a generalization of DPO that supports multiple preference dimensions and dynamic interpolation via a simplex-weighted $\lambda$ vector. Our method enables listwise supervision and flexible alignment without re-training. While our experiments are conducted on 1B–2B scale models, this is an intentional choice: smaller models provide a more stringent testbed where performance improvements more clearly reflect the effectiveness of the alignment strategy itself. Moreover, such models are widely used in compute-constrained applications, making our improvements both methodologically meaningful and practically valuable. Empirical results show that our approach matches or surpasses standard DPO on alignment benchmarks while offering improved adaptability.
@@ Introduction
Large language models (LLMs) have achieved impressive results across natural language tasks, but often misalign with human preferences—producing responses that are factually incorrect, verbose, or unhelpful. This misalignment stems from the next-token prediction objective, which lacks explicit human feedback.
To address this, recent approaches incorporate preference supervision. RLHF trains a reward model from human comparisons and optimizes it via reinforcement learning, but suffers from high cost, instability, and deployment complexity. Direct Preference Optimization (DPO) offers a simpler alternative by directly optimizing model outputs to reflect pairwise preferences, achieving strong results with lower training overhead. However, DPO assumes a single reward dimension and cannot adapt to varying or multi-faceted user intents.
In practice, alignment involves multiple criteria—e.g., helpfulness, honesty, fluency—whose importance may shift across tasks. To address this, we propose Multi-Preference Lambda-weighted Listwise DPO, a generalization of DPO that incorporates multiple preference axes and allows dynamic control via a weight vector $\lambda$ drawn from the simplex.
Our method extends DPO from pairwise to listwise supervision and introduces a learnable scheduling mechanism over $\lambda$ vectors, enabling flexible, fine-grained alignment. Empirical results on multi-objective benchmarks show that our method matches or exceeds standard DPO while offering greater control and adaptability.
@@
Reinforcement learning from human feedback (RLHF) [18, 1] aligns LLMs via reward models and PPO [21], but is costly and unstable [28, 5]. Direct Preference Optimization (DPO) [19] simplifies alignment by reframing it as binary classification on preference pairs, improving stability and efficiency. Extensions like Quark [26] and LLaMA-DPO [16] show strong performance, though limited to fixed, pairwise objectives.
@@
Listwise learning captures global ranking better than pairwise [4, 24]. UltraFeedback [8] provides ranked human feedback, and methods like GLOW [25] train reward models from such data. Our approach differs by optimizing listwise loss directly in policy space, without intermediate reward modeling.
@@
Prior work on multi-reward RL [12], decomposition [13], and reward interpolation [23] requires retraining for each preference trade-off. Our method incorporates a weight vector $\lambda$ during training, enabling generalization to arbitrary preferences. Related ideas include universal value function approximators [20]. Steerable LMs using prompts [3, 22], control tokens [11], or constrained decoding [17] offer limited flexibility. In contrast, our method enables fine-grained, multi-dimensional alignment via $\lambda$ without modifying prompts or decoding.
@@
Let $x$ denote the prompt, and $y_{w}\succ y_{l}$ represent a human preference between a pair of completions: $y_{w}$ (the preferred answer) and $y_{l}$ (the less preferred one). These preferences are sampled from an underlying distribution $p^{*}(y_{w},y_{l}|x)$, which is typically assumed to be induced by a latent reward function $r(x,y)$. A commonly adopted model is the Bradley-Terry (BT) model [2]:
@@
$\mathcal{L}_{\text{BT}}(r)=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma(r(x,y_{w})-r(x,y_{l}))\right],$ (2)
@@
$\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(y|x)}\left[r(x,y)-\beta\cdot\mathrm{KL}\left(\pi_{\theta}(y|x)\|\pi_{\text{ref}}(y|x)\right)\right],$ (3)
@@
This objective encourages the policy to produce high-reward outputs while staying close to $\pi_{\text{ref}}$. However, in practice, it requires sampling, reward evaluation, and reinforcement learning algorithms such as PPO [21] to optimize.
@@
$\pi_{r}(y|x)\propto\pi_{\text{ref}}(y|x)\cdot\exp\left(\frac{1}{\beta}r(x,y)\right).$ (4)
@@
$\beta(r(x,y_{w})-r(x,y_{l}))=\log\frac{\pi_{r}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{r}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}.$ (5)
@@
$\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right].$ (6)
@@
Motivated by the limitations of pairwise DPO under a single fixed preference distribution, we introduce a generalized framework that supports listwise preference supervision across multiple human objectives. In real-world settings, human judgments often reflect diverse criteria—such as helpfulness, harmlessness, and conciseness—and capturing these via a unified formulation enables controllable alignment across axes of interest.
Our key insight is that we can reparameterize multi-dimensional listwise preferences using a convex combination of multiple reward signals. Specifically, we define a simplex-weighted listwise loss that extends the DPO framework to dynamic, multi-objective optimization.
@@
$P_{\theta}(y_{i}|x)=\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}},$ (7)
@@
$\mathcal{L}_{\text{listwise}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[\sum_{i=1}^{N}p^{*}(y_{i}|x)\cdot\log P_{\theta}(y_{i}|x)\right].$ (8)
@@
$p^{\lambda}(y_{i}|x)=\sum_{k=1}^{m}\lambda_{k}\cdot p^{*(k)}(y_{i}|x).$ (9)
@@
$\mathcal{L}_{\lambda\text{-DPO}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}|x)\cdot\log P_{\theta}(y_{i}|x)\right].$ (10)
@@
$\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}=-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}|x)-P_{\theta}(y_{i}|x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)\right].$ (11)
@@
In practice, we can reuse datasets like UltraFeedback [8], where multiple preference criteria are annotated. During inference, changing $\lambda$ enables personalized alignment without modifying model parameters.
@@
While Direct Preference Optimization (DPO) offers a robust framework for aligning language models with human preferences via pairwise supervision, it inherently simplifies the preference space by assuming binary feedback between two completions. In this section, we analyze the theoretical implications of this simplification and motivate the use of a listwise and multi-preference extension.
@@
Given a prompt $x$ and two outputs $y_{w}$ (winner) and $y_{l}$ (loser), DPO defines a binary preference distribution:
$P^{*}(y_{i}|x)=\begin{cases}1&\text{if }y_{i}=y_{w}\\ 0&\text{if }y_{i}=y_{l}\end{cases}$ (12)
This implicitly constrains the learning objective to satisfy $P^{*}(y_{w}|x)+P^{*}(y_{l}|x)=1$, ignoring the quality of other plausible candidates. Consequently, the optimized policy $\pi_{\theta}$ may fail to capture the full spectrum of user preferences, introducing estimation error.
Listwise Distribution via Softmax.
To mitigate this, we define a listwise target distribution $\pi^{*}$ over all completions $\{y_{1},y_{2},\dots,y_{N}\}$ for prompt $x$, derived from scalar reward scores $r(x,y_{i})$:
@@
This distribution captures relative preference over all candidates, allowing for more representative updates and lower variance.
Multi-Preference Supervision.
Real-world alignment involves multiple competing objectives, such as helpfulness, safety, and factuality. Let $p^{*(k)}(y_{i}|x)$ denote the listwise preference under objective $k\in\{1,\dots,m\}$. We define a convex combination of these objectives via a preference vector $\lambda\in\Delta^{m}$:
@@
The training objective minimizes the cross-entropy between the aggregated distribution $p^{\lambda}$ and the policy-induced distribution $P_{\theta}$:
@@
Gradient Interpretation.
The gradient of this objective is:
$\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}=-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}\mid x)-P_{\theta}(y_{i}\mid x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}\mid x)\right]$ (16)
This update rule drives $\pi_{\theta}$ toward the weighted human preference distribution $p^{\lambda}$, allowing alignment to reflect fine-grained and multi-dimensional values.
Variance Reduction via Listwise Supervision.
Using a normalized softmax distribution $\pi^{*}$ derived from reward scores avoids the instability and high variance that arise when optimizing based on hard binary labels. Empirically, listwise training improves sample efficiency and convergence by utilizing the full output ranking structure.
@@
In summary, Multi-Preference Lambda-weighted Listwise DPO:
  • Recovers a more informative supervision signal via $\pi^{*}$.
  • Reduces estimation error and variance in gradient updates.
  • Enables alignment over a continuous space of user-defined objectives via $\lambda$.
Experiment Setup
Experimental Objectives and Comparison Targets
We consider two levels of comparison:
Model comparison: We experiment with two strong open-source language models—LLaMA and Qwen. These models differ in architecture, pretraining corpora, and tokenizer, providing a robust testbed for evaluating generalizability.
Method comparison: We compare DPO and $\lambda$-DPO. For each method and model pair, we perform independent hyperparameter tuning (e.g., learning rate, batch size, $\beta$) to ensure that comparisons are fair and reflect the best achievable performance under comparable resource constraints.
Training Settings
Inspired by the SimPO baseline setup, we adopt two training pipelines for both models:
  • Base: The model is directly trained on the UltraFeedback dataset using DPO or $\lambda$-DPO. No supervised fine-tuning is applied before preference optimization. This setting isolates the influence of the preference learning method alone.
  • Finetuned-Base: The model is first supervised fine-tuned (SFT) on the UltraChat-200k dataset, which contains approximately 200K high-quality dialogue examples. The resulting SFT model is then optimized using UltraFeedback preferences via DPO or $\lambda$-DPO. This pipeline reflects realistic alignment procedures.
All training is conducted on A100 GPUs, with consistent batch sizes and sequence lengths. For each setting, hyperparameters are tuned via validation performance on reward preference accuracy and KL stability.
Evaluation Benchmarks and Metrics
We evaluate model performance using the lm-eval-harness [9] framework, which tests accuracy on a diverse set of benchmarks covering reasoning, factual correctness, and commonsense understanding.
  • MMLU [10]: 57 subject areas with around 15K test questions targeting academic and professional knowledge.
  • ARC (Challenge) [6]: Grade-school science questions ($\sim$1.1K) requiring multi-step reasoning.
  • HellaSwag [27]: 10K commonsense sentence completion tasks with adversarial distractors.
  • TruthfulQA [15]: 817 questions designed to expose hallucinations and factual errors.
  • Winograd [14]: 273 examples testing pronoun coreference and syntactic disambiguation.
  • GSM8K [7]: 8.5K arithmetic reasoning problems with step-by-step solutions.
Ablation Experiment
[…]
  • Unimodal $\lambda$: we use fixed one-hot vectors (e.g., $(1,0,0,\dots)$), effectively disabling interpolation and isolating optimization to a single preference dimension per instance. This variant serves as a control version without multi-dimensional aggregation.
This comparison helps isolate the contribution of $\lambda$-based preference interpolation within the listwise setting. For a separate comparison between standard DPO and our listwise formulation, please refer to the main benchmark results in Table 1.
@@
We compare standard DPO and our proposed $\lambda$-DPO under identical training and evaluation setups on two models: LLaMA 3.2B Instruct and Qwen 1.8B Chat. Both models are first supervised fine-tuned (SFT) on UltraChat-200k, then aligned using the UltraFeedback listwise dataset.
@@
Evaluation.
We evaluate on six benchmarks (MMLU, ARC, HellaSwag, TruthfulQA, WSC273, GSM8K) using the lm-eval-harness, reporting accuracy as the primary metric.
@@
Table 1 reports the accuracy of DPO and $\lambda$-DPO on six representative benchmarks for both LLaMA 3.2B Instruct and Qwen 1.8B Chat. On LLaMA, $\lambda$-DPO consistently improves performance across all tasks, with notable gains such as +1.1% on ARC, +3.3% on GSM8K, and +3.6% on TruthfulQA. The improvement on WSC273 is also evident (+1.8%). For Qwen 1.8B Chat, $\lambda$-DPO outperforms DPO on all benchmarks except TruthfulQA, where DPO achieves a slightly higher score. Notably, large gains are observed on GSM8K (+14.7%) and MMLU (+3.9%), demonstrating the generalization benefits of $\lambda$-DPO across different model families.
@@
LLaMA 3.2B Instruct DPO 0.458 0.364 0.456 0.378 0.696 0.312
@@
Qwen 1.8B Chat DPO 0.379 0.271 0.418 0.366 0.656 0.008
@@
Table 1: Accuracy comparison between DPO and $\lambda$-DPO on six benchmarks using LLaMA 3.2B Instruct and Qwen 1.8B Chat. Bold values indicate better performance for each benchmark.
@@
Nonetheless, for certain adversarial or logic-heavy tasks like TruthfulQA, the benefits of listwise supervision may be less reliable. Integrating $\lambda$-DPO with structured supervision—such as chain-of-thought prompting [22]—could potentially mitigate this gap.
@@
In summary, $\lambda$-DPO demonstrates consistent improvements over DPO in most evaluation scenarios, confirming the effectiveness of listwise preference supervision. Its generalization advantage makes it a strong candidate for broader alignment tasks. Future work may explore hybrid strategies that combine $\lambda$-DPO with structured reasoning methods to further enhance performance on complex benchmarks.
@@
We observe that all four models achieve relatively comparable performance across the six benchmarks. The $\lambda=[0,1,0,0]$ model (favoring honesty) achieves the highest average accuracy, while the $\lambda=[1,0,0,0]$ model (favoring helpfulness) performs best on truthfulqa and shows strong performance on gsm8k. The models focusing on instruction-following and fluency show only marginally lower scores, with no major degradation in benchmark performance.
@@
In future work, we plan to extend this ablation framework by exploring:
  • Adaptive $\lambda$ scheduling, where preference weights change based on input context or task type.
  • Hierarchical modeling that separates semantic utility from stylistic fluency.
  • Persona- or task-specific interpolation strategies to match different downstream alignment needs.

### v5

# Multi-Preference Lambda-weighted Listwise DPO for Small-Scale Model Alignment

@@ Abstract
Large language models (LLMs) demonstrate strong generalization across a wide range of language tasks, but often generate outputs that misalign with human preferences. Reinforcement Learning from Human Feedback (RLHF) addresses this by optimizing models toward human preferences using a learned reward function and reinforcement learning, yielding improved alignment but suffering from high computational cost and instability. Direct Preference Optimization (DPO) simplifies the process by treating alignment as a classification task over binary preference pairs, reducing training overhead while achieving competitive performance. However, it assumes fixed, single-dimensional preferences and only supports pairwise supervision. To address these limitations, we propose Multi-Preference Lambda-weighted Listwise DPO, which allows the model to learn from more detailed human feedback and flexibly balance multiple goals such as helpfulness, honesty, and fluency. Our method models full-ranked preference distributions rather than binary comparisons, enabling more informative learning signals. The $\lambda$ vector controls the relative importance of different alignment goals, allowing the model to generalize across diverse human objectives. During inference, $\lambda$ can be adjusted without retraining, providing controllable alignment behavior for downstream use. We also introduce a learned scheduler that dynamically samples performant $\lambda$ configurations to improve robustness. Notably, our method requires only 20GB of GPU memory for training, making it suitable for compute-constrained settings such as academic labs, educational tools, or on-device assistants. Experiments on 1B–2B scale models show that our method consistently outperforms standard DPO on alignment benchmarks while enabling efficient, controllable, and fine-grained adaptation suitable for real-world deployment.
@@ Introduction
Large language models (LLMs) such as GPT (Brown et al. 2020) and LLaMA (Touvron and et al. 2023) have emerged as powerful backbones in natural language processing, demonstrating impressive capabilities across tasks such as question answering, summarization, and instruction following. However, despite their success, LLMs often produce outputs that do not align well with human expectations and values.
This misalignment frequently manifests as responses that are verbose, inaccurate, or unhelpful. It stems from their pretraining objective: next-token prediction, which emphasizes statistical co-occurrence over communicative intent, and inherently lacks exposure to explicit human feedback (Ouyang et al. 2022).
To address this, a line of alignment techniques has been developed to incorporate preference supervision. Among them, Reinforcement Learning from Human Feedback (RLHF) has become a standard paradigm (Christiano et al. 2017; Ouyang et al. 2022; Bai et al. 2022). By training a reward model from human comparisons and optimizing it via reinforcement learning (e.g., PPO (Schulman et al. 2017)), RLHF enables models to better match human values. However, the reliance on expensive reward modeling, unstable optimization dynamics, and high computational cost make RLHF difficult to scale and deploy reliably (Ziegler et al. 2019).
As a lightweight alternative, Direct Preference Optimization (DPO) has recently been proposed (Rafailov et al. 2023). DPO reframes alignment as a classification problem over pairwise preferences, directly optimizing model outputs to prefer preferred responses over less preferred ones. Without requiring an explicit reward model or reinforcement loop, DPO achieves strong alignment with reduced training overhead. Nevertheless, existing DPO methods typically assume a single, fixed alignment objective and binary supervision. This limits their applicability in practical scenarios, where alignment is inherently multi-dimensional—balancing objectives such as helpfulness, factuality, honesty, and fluency (Bai et al. 2022; Wu et al. 2021).
To accommodate this diversity, several controllable generation strategies have been explored, such as prompt tuning (Liu and et al. 2021), control tokens (Keskar et al. 2019), and multi-reward reinforcement learning (Lee et al. 2023). However, these approaches often require retraining under each new preference configuration, introduce architectural complexity, or generalize poorly in unseen alignment regimes.
In this work, we propose Multi-Preference Lambda-weighted Listwise DPO, a unified framework that extends the standard DPO formulation along two critical dimensions: supervision granularity and preference flexibility. Concretely, we generalize pairwise supervision to listwise preference modeling, enabling the model to learn from multiple ranked completions per prompt. This richer supervision signal allows the model to reason over global ranking structures, rather than relying on isolated binary comparisons. As a result, it improves gradient quality and better captures nuanced distinctions between high-quality and suboptimal responses. Simultaneously, we introduce a lambda-weighted composition mechanism, in which a preference vector $\lambda\in\Delta^{m}$ governs the relative importance of each alignment axis. This design allows a single model to flexibly interpolate between multiple objectives without retraining.
To further enhance adaptability, we develop a learned scheduling strategy that dynamically adjusts the sampling distribution over $\lambda$ vectors based on their empirical performance.
Empirical results on two strong 1B–2B scale LLMs (LLaMA and Qwen) across six alignment benchmarks demonstrate that our approach consistently outperforms standard DPO. Furthermore, by adjusting $\lambda$ during inference, our method supports personalized alignment without modifying model weights, making it highly practical for real-world deployment in education, conversational agents, and lightweight on-device applications.
@@
Reinforcement learning from human feedback (RLHF) (Ouyang et al. 2022; Bai et al. 2022) has become a widely used paradigm for aligning large language models (LLMs) with human preferences. It typically involves training a reward model using pairwise human comparisons and then optimizing the base language model using reinforcement learning algorithms such as Proximal Policy Optimization (PPO) (Schulman et al. 2017). This two-stage approach has demonstrated strong alignment performance, notably in systems like InstructGPT (Ouyang et al. 2022) and Claude (Bai et al. 2022). However, RLHF suffers from several limitations: reward models can be noisy or biased, PPO introduces high variance and instability during training, and the overall pipeline is computationally intensive and hard to scale (Ziegler et al. 2019; Christiano et al. 2017).
To reduce this complexity, direct preference optimization (DPO) (Rafailov et al. 2023) reframes preference alignment as a supervised classification problem. Instead of learning a reward model, DPO directly optimizes the policy to prefer preferred responses using a KL-regularized binary cross-entropy loss. This improves both training stability and implementation simplicity. Variants such as LLaMA-DPO (Liu et al. 2023) and Quark (Yao et al. 2023) have extended this framework with architectural tuning and reward quantization, showing strong performance in single-objective settings. However, most existing work in this area assumes fixed binary preferences and lacks the flexibility to accommodate multiple alignment goals or dynamically varying human objectives.
@@
While pairwise preference modeling is common, listwise preference learning has been shown to better capture global ranking information (Burges et al. 2005; Xia et al. 2008). Instead of treating comparisons independently, listwise approaches model the entire ranking over a set of candidates, providing richer supervision and reducing sample complexity. In the context of language model alignment, UltraFeedback (Du et al. 2023) provides large-scale human feedback data with complete ranking annotations over multiple completions. This enables more expressive and fine-grained alignment signals beyond binary feedback.
Recent works such as GLOW (Xu et al. 2023) leverage such ranked data to train reward models using listwise objectives. These models are then used in RLHF-style pipelines to align policies. However, they still rely on intermediate reward learning and reinforcement optimization. In contrast, our method directly optimizes the policy using listwise cross-entropy loss, bypassing the need for explicit reward modeling. This not only simplifies the training pipeline but also provides more stable gradients and better captures the structure of human preferences.
@@
Human alignment is inherently multi-faceted, involving trade-offs among objectives such as helpfulness, honesty, informativeness, and harmlessness. Prior work has explored multi-objective reinforcement learning frameworks (Korbak et al. 2023; Lee et al. 2023) that combine multiple reward signals through scalarization or decomposition. However, these methods typically require retraining or manual tuning of trade-off weights for each new configuration. Similarly, reward interpolation techniques (Wu et al. 2021) attempt to blend preferences using convex combinations of pretrained reward models, but they still involve separate optimization for each alignment setting.
To enable better generalization, our method integrates a weight vector $\lambda$ into the training process, allowing the model to internalize a continuous spectrum of preference trade-offs. This is conceptually related to universal value function approximators (Schaul et al. 2015), which generalize value functions across goals or reward vectors. In the alignment context, we treat $\lambda$ as a parameter controlling the target alignment behavior, enabling downstream control without model retraining.
Alternative strategies for controllable generation include prompt-based steering (Brown et al. 2020; Wei et al. 2022), control tokens (Keskar et al. 2019), and constrained decoding techniques (Lu et al. 2022). While effective in certain settings, these methods often struggle with consistency, lack semantic grounding in human-aligned dimensions, or require careful hand-engineering. Our approach differs in that it integrates multi-dimensional alignment directly into the learning objective, enabling fine-grained, interpretable control through $\lambda$ without modifying prompts, decoding strategies, or model architecture.
@@
Let $x$ denote the prompt, and $y_{w}\succ y_{l}$ represent a human preference between a pair of completions: $y_{w}$ (the preferred answer) and $y_{l}$ (the less preferred one). These preferences are sampled from an underlying distribution $p^{*}(y_{w},y_{l}|x)$, which is typically assumed to be induced by a latent reward function $r(x,y)$. A commonly adopted model is the Bradley-Terry (BT) model (Bradley and Terry 1952):
@@
$\mathcal{L}_{\text{BT}}(r)=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma(r(x,y_{w})-r(x,y_{l}))\right]$ (2)
@@
$\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(y|x)}\left[r(x,y)-\beta\cdot\mathrm{KL}\left(\pi_{\theta}(y|x)\|\pi_{\text{ref}}(y|x)\right)\right]$ (3)
@@
This objective encourages the policy to produce high-reward outputs while staying close to $\pi_{\text{ref}}$. However, in practice, it requires sampling, reward evaluation, and reinforcement learning algorithms such as PPO (Schulman et al. 2017) to optimize.
@@
$\pi_{r}(y|x)\propto\pi_{\text{ref}}(y|x)\cdot\exp\left(\frac{1}{\beta}r(x,y)\right)$ (4)
@@
$\beta(r(x,y_{w})-r(x,y_{l}))=\log\frac{\pi_{r}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{r}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}$ (5)
@@
$-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right]$ (6)
@@
Figure 1: Overview of the $\lambda$-DPO architecture. The system integrates listwise preference modeling and a controllable $\lambda$ vector to align model outputs with multi-dimensional human preferences.
@@
Motivated by the limitations of pairwise DPO under a single fixed preference distribution, we propose two orthogonal enhancements: listwise supervision and simplex-weighted optimization. First, instead of relying on binary comparisons, we adopt listwise learning to utilize full human rankings over multiple outputs. This provides stronger supervision signals and captures richer preference structure. Second, we generalize DPO to support multi-objective alignment by introducing a simplex-weighted aggregation over multiple reward components. The resulting $\lambda$-DPO formulation enables flexible interpolation across alignment dimensions such as helpfulness, harmlessness, and fluency. Together, these extensions allow controllable and scalable alignment with more nuanced human feedback.
@@
$P_{\theta}(y_{i}|x)=\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}}$ (7)
@@
$\mathcal{L}_{\text{listwise}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[\sum_{i=1}^{N}p^{*}(y_{i}|x)\cdot\log P_{\theta}(y_{i}|x)\right]$ (8)
@@
$p^{\lambda}(y_{i}|x)=\sum_{k=1}^{m}\lambda_{k}\cdot p^{*(k)}(y_{i}|x)$ (9)
@@
$\mathcal{L}_{\lambda\text{-DPO}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}|x)\cdot\log P_{\theta}(y_{i}|x)\right]$ (10)
@@
$\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}=-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}|x)-P_{\theta}(y_{i}|x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)\right]$ (11)
@@
In practice, we can reuse datasets like UltraFeedback (Du et al. 2023), where multiple preference criteria are annotated. During inference, changing $\lambda$ enables personalized alignment without modifying model parameters.
@@
While Direct Preference Optimization (DPO) offers a principled and scalable approach to aligning language models with human feedback, it is fundamentally limited by its reliance on pairwise supervision. In this section, we provide a detailed theoretical analysis of these limitations and motivate our proposed extensions—namely, listwise supervision and preference mixing via simplex weighting—to overcome them in multi-objective alignment settings.
@@
Given a prompt $x$ and a pair of sampled completions $y_{w}$ (preferred) and $y_{l}$ (dispreferred), DPO defines a binary target distribution over the pair as:
$p^{*}(y_{i}\mid x)=\begin{cases}1&\text{if }y_{i}=y_{w}\\ 0&\text{if }y_{i}=y_{l}\end{cases}$ (12)
This construction is simple and efficient, allowing the loss to be computed from binary human preferences. However, it treats the preference signal as a hard decision, discarding nuanced feedback such as “somewhat better” or “equally good.” Furthermore, the normalization over only two samples induces high-variance gradients, as the updates depend on stochastic comparisons between limited completions. Since only the relative preference between $y_{w}$ and $y_{l}$ is modeled, all other completions (which might be equally or more suitable) are ignored. This can lead to suboptimal generalization, especially when the number of plausible high-quality completions is large.
Listwise Distribution and Multi-Preference Supervision.
To address the expressiveness gap, we generalize DPO to operate over a list of completions $\{y_{1},\dots,y_{N}\}$ using listwise supervision. Specifically, given a reward model $r(x,y_{i})$ that quantifies the quality of each completion, we define a normalized target distribution over the list using softmax:
@@
This formulation assigns probability mass proportionally to reward values, providing smoother gradients and improving credit assignment. Unlike the binary setting, listwise DPO makes use of the full ranking structure, offering a richer training signal and faster convergence.
Moreover, human preferences in practical applications are inherently multi-dimensional—comprising objectives like helpfulness, safety, factuality, politeness, and conciseness. Suppose we define $m$ distinct reward models, each representing one objective. We can construct a family of listwise distributions $\{p^{*(k)}(y_{i}|x)\}_{k=1}^{m}$, each aligned with a single criterion. A convex combination of these objectives is parameterized by a preference vector $\lambda\in\Delta^{m}$, forming a new composite distribution:
@@
This formulation enables controllable alignment by adjusting $\lambda$, allowing practitioners to emphasize certain values over others (e.g., prioritize safety over informativeness). The model is trained to match this target distribution via cross-entropy:
@@
Its gradient reveals the direction of improvement, given by $\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}=$
$-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}\mid x)-P_{\theta}(y_{i}\mid x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}\mid x)\right]$ (16)
which encourages the model’s policy distribution to match the aggregated preference signal from multiple human value dimensions.
@@
In summary, Multi-Preference Lambda-weighted Listwise DPO offers a theoretically grounded generalization of the original DPO framework. First, it extends binary supervision to listwise distributions, providing smoother and more informative gradient updates. Second, it integrates diverse human objectives through a simplex-weighted mixture of reward-aligned targets, enabling dynamic and tunable preference alignment. Finally, it reduces sample inefficiency and training instability by leveraging full output rankings and avoiding overfitting to binary labels, thus making it well-suited for real-world alignment tasks with conflicting value axes.
@@
We compare standard DPO and our proposed $\lambda$-DPO under identical training and evaluation setups on two models: LLaMA 3.2-1B Instruct and Qwen 1.8B Chat. Both models are first supervised fine-tuned (SFT) on UltraChat-200k, then aligned using the UltraFeedback listwise dataset.
@@
Table 1 reports the accuracy of DPO and $\lambda$-DPO on six representative benchmarks for both LLaMA 3.2-1B Instruct and Qwen 1.8B Chat. On LLaMA, $\lambda$-DPO consistently improves performance across all tasks, with notable gains such as +1.1% on ARC, +3.3% on GSM8K, and +3.6% on TruthfulQA. The improvement on WSC273 is also evident (+1.8%). For Qwen 1.8B Chat, $\lambda$-DPO outperforms DPO on all benchmarks except TruthfulQA, where DPO achieves a slightly higher score. Notably, large gains are observed on GSM8K (+14.7%) and MMLU (+3.9%), demonstrating the generalization benefits of $\lambda$-DPO across different model families.
@@
LLaMA 3.2-1B Instruct ORPO 0.456 0.355 0.447 0.435 0.678 0.327
SIMPO 0.459 0.362 0.455 0.381 0.703 0.318
DPO 0.458 0.364 0.456 0.378 0.696 0.312
@@
Qwen 1.8B Chat ORPO 0.377 0.278 0.417 0.347 0.661 0.007
SIMPO 0.415 0.295 0.423 0.355 0.658 0.007
DPO 0.379 0.271 0.418 0.366 0.656 0.008
@@
Table 1: Accuracy comparison across six benchmarks. The first block compares ORPO, SIMPO, DPO, and $\lambda$-DPO on LLaMA 3.2-1B Instruct. The second block compares the same four methods on Qwen 1.8B Chat. Bold values indicate the best performance within each model group.
@@
Nonetheless, for certain adversarial or logic-heavy tasks like TruthfulQA, the benefits of listwise supervision may be less reliable. Integrating $\lambda$-DPO with structured supervision—such as chain-of-thought prompting (Wei et al. 2022)—could potentially mitigate this gap.
@@
In summary, $\lambda$-DPO brings two key improvements over standard DPO. First, by adopting listwise preference supervision, it leverages full ranking information instead of pairwise comparisons, resulting in richer and lower-variance gradient signals. Second, through the introduction of the $\lambda$ vector, it supports flexible interpolation across multiple alignment objectives (e.g., helpfulness, harmlessness, factuality), enabling controllable alignment behavior and improved generalization under diverse user preferences. These two enhancements jointly contribute to the observed performance gains across benchmarks. Future work may explore hybrid strategies that combine $\lambda$-DPO with structured reasoning techniques to further boost alignment quality on complex tasks.
@@
All four models perform similarly across benchmarks. The honesty-focused model ($\lambda=[0,1,0,0]$) achieves the highest average accuracy, while the helpfulness-focused model ($\lambda=[1,0,0,0]$) excels on truthfulqa and performs well on gsm8k. Other preference models show only slight drops, with no significant performance degradation.